Income Statement and balance sheet information abstracted from a recent annual report of Wolverine World Wide, Inc., appears below:

Balance Sheet (in millions):

December 30, 2017 December 31, 2016
Current Assets: Inventories $276.7 $348.7

Income Statement (in millions):

For the year ended December 30,2017 For the year ended December 31, 2016
Net Sales $2350.0 $2494.6
Cost of Goods Sold $1426.6 $1526.4
Gross Profit $923.4 $968.2

The significant accounting policies note disclosure contained the following:

Inventories

The Company used the LIFO method to value inventories of $53.2 million at December 30, 2017 and $66.2 million at December 31, 2016. During fiscal year 2017, a reduction in inventory quantities resulted in a liquidation of applicable LIFO inventory quantities carried at lower costs in previous years. This LIFO liquidation decreased cost of goods sold by $6.0 million. If the FIFO method had been used, inventories would have been $16.4 million $22.4 million higher reported at December 30, 2017 and December 31, 2016 respectively.

2. Calculate what beginning inventory and ending inventory would have been for the year ended December 31, 2017 if Wolverine had used FIFO for all of its inventories.

3. Calculate what cost of goods sold and profit would have been for the year ended December 31, 2017, if Wolverine had used FIFO for all its inventories.

  (For all requirements, do not round intermediate calculations. Enter your answers in millions rounded to 1 decimal place.)

2-i.

Beginning inventory million
2-ii. Ending inventory    million
3-i. Cost of Goods Sold million
3-ii. Gross Profit million
Answer #1
2 Ending inventory Beginning inventory
Inventories as per balance sheet 276.7 348.7
Increase due to change from LIFO to FIFO 16.4 22.4
FIFO inventory balance 293.1 371.1
3 Increase in ending inventory=$ 16.4 million
Increase in beginning inventory=$ 22.4 million
Net increase in beginning inventory=22.4-16.4=$ 6 million
Increase in beginning inventory will result in increase in cost of goods sold by $ 6 million
Cost of goods sold=1426.6+6=$ 1432.6 million
Profit=Net sales-Cost of goods sold=2350-1432.6=$ 917.4
Summary:
2-i. Beginning inventory 371.1 Million
2-ii. Ending inventory 293.1 Million
3-i. Cost of goods sold 1432.6 Million
3-ii. Gross profit 917.4 Million